Find the smallest value of P=|z-2|^2 + |z+1-i|^2 + | z-2 -5i|
And z (z = x +yi : x,y are real numbers) is a complex numbers satisfies the condition 2|(x+yi)-1-2i| = |3i + 1 - 2(x-yi)|
1
Expert's answer
2021-07-13T09:24:44-0400

Let us develop the expression of "P=|z-2|^2+|z+1-i|^2+|z-2-5i|" in terms of "x" and "y" :

"P=(x-2)^2+y^2+(x+1)^2+(y-1)^2+\sqrt{(x-2)^2+(y-5)^2}"

And the condition is "4(x-1)^2+4(y-2)^2=(1-2x)^2+(3+2y)^2".

The condition can be simplified (after developing all the squares) as

"-4x-28y+10=0"

or "2x+14y=5"

Now as this condition admits a reformulation in the form "x=7y-2.5", we can substitue this into the expression of "P" :

"P=(7y-4.5)^2+y^2+(7y-1.5)^2+(y-1)^2+\sqrt{(7y-4.5)^2+(y-5)^2}"

Which we can rewrite as

"P=100y^2-86y+23.5+\sqrt{50y^2-73y+45.25}"

Now by calculating "P'=200y-86+\frac{100y-73}{2\sqrt{50y^2-73y+45.25}}"

By finding the root and submitting it into the expression of "P" yields us that the minimum of "P" in the given conditions is "P(x_{min},y_{min})\approx 9.7934"